# Revisiting Non-Rotating Star Models: Classical Existence and Uniqueness Theory and Scaling Relations

### v1

@@ 2.2 Existence Results for Non-rotating Bodies
For Part (vii), we already know $\lambda_{m}<0$. Moreover, (18) (19), together with $\sigma_{m}$ is continuous, give (EL), see for example [16, Section 5] or [7, Theorem 2.25]. (Part (vii))
@@
For Part (ii), we only consider the case that complies with polytropic equations of state $P(\sigma)=K\sigma^{\gamma}$. The general case can be found in [13, Theorem 3 (d)] [^4] However, it should be noted that the argument there is set within a quantum mechanical framework and does not include an internal energy term, although McCann mentioned that the results can be applied “equally well to all $A(\rho)$ consistent with (F1)-(F3)[16, Section 3].. We utilize the scaling relations mentioned in Section 3. That is, given $\sigma$ a non-rotating minimizer with mass $1$, then $\sigma_{m}(x)=\frac{1}{A}\sigma\left(\frac{1}{B}x\right)$ will be a non-rotating minimizer with mass $m$, where $A(m)=m^{-\frac{2}{3\gamma-4}},B(m)=m^{\frac{\gamma-2}{3\gamma-4}}$. At the same time, we know $e_{0}(m)=E_{0}\left(\sigma_{m}\right)=E_{0}\left(\frac{1}{A}\sigma\left(\frac{1}{B}x\right)\right)=m^{\frac{5\gamma-6}{3\gamma-4}}e_{0}$. Notice when $m=0$, such expression also holds since $e_{0}(0)=0$. By the similar arguments as the proof of Step (2) (Lemma 2.25), we know $e_{0}\coloneq e_{0}(1)<0$, which implies $e_{0}(m)$ decreases continuously from $e_{0}(0)=0$ and is strictly concave (Part (ii)).
@@
For Part (viii), first the concavity of $e_{0}(m)$ implies the existence of left-hand and righthand derivations (since they are monotonic), and they are equal a.e. (since concavity implies Lipschitz continuity and then apply Rademacher’s theorem [10, Subsection 5.8.3]). Thanks to Euler-Lagrange equations (EL), we multiply $\sigma_{m}$ on both sides and integrate, and then we obtain $\int E_{0}^{\prime}\left(\sigma_{m}\right)\sigma_{m}=\lambda_{m}\cdot m$. Notice due to Part (iv) and (v), we know $\sigma_{m}\in P_{\infty}\left(\sigma_{m}\right)$ thus $\int E_{0}^{\prime}\left(\sigma_{m}\right)\sigma_{m}$ makes sense and $\int E_{0}^{\prime}\left(\sigma_{m}\right)\sigma_{m}=\lim\limits_{t\rightarrow 0^{+}}\frac{E_{0}\left(\sigma_{m}+t\sigma_{m}\right)-E_{0}\left(\sigma_{m}\right)}{t}$ $\lim\limits_{t\rightarrow 0^{-}}\frac{E_{0}\left(\sigma_{m}\right)-E_{0}\left(\sigma_{m}-t\sigma_{m}\right)}{t}$. Since $\sigma_{m}+t\sigma_{m}$ has mass $(1+t)m,e_{0}((1+t)m)=E_{0}\left(\sigma_{(1+t)m}\right)\leq E_{0}\left(\sigma_{m}+t\sigma_{m}\right)$, thus $\lim\limits_{t\rightarrow 0^{+}}\frac{E_{0}\left(\sigma_{m}+t\sigma_{m}\right)-E_{0}\left(\sigma_{m}\right)}{t}\geq\lim\limits_{t\rightarrow 0^{+}}\frac{e_{0}((1+t)m)-e_{0}(m)}{t}=me_{0}^{\prime}\left(m^{+}\right)$, then we have $\lambda_{m}\geq e_{0}^{\prime}\left(m^{+}\right)$. Similarly, we have $\lambda_{m}\leq e_{0}^{\prime}\left(m^{-}\right)$. (Part (viii))
@@
For Part (ix), thanks to Part (vi), we know the global continuity of $\sigma_{m}$, and then we can apply the same arguments as in the proof of [7, Theorem 2.25 (vii)] and show the existence of $\nabla P\left(\sigma_{m}\right)$ on the boundary $\partial\left\{\sigma_{m}>0\right\}$, and verify $\sigma_{m}$ is a solution to reduced Euler-Poisson equations (EP’) in $\mathbb{R}^{3}$. (Part (ix)) ∎
@@
When we prove Theorem 2.6 (vi), we mention $\sigma_{m}{}^{\frac{1}{3}}\leq cV_{\sigma_{m}}$ a.e. on the set where $\sigma_{m}\geq K$ for some $K$ large enough. Actually it turns out even if we do not know $\lambda_{m}<0$, $\sigma_{m}{}^{\frac{1}{3}}\leq cV_{\sigma_{m}}$ can still holds true since we can absorb $\lambda_{m}$ when $\sigma_{m}$ is large enough. See [1, Lemma 3].
@@
We point out in Theorem 2.6 we can show the existence of non-rotating single star no matter how large the mass is. However, if we replace (F3) by (F3’), then we can only obtain minimizers with total mass not larger than $m(K)$ via variational method for some $m(K)>0$. $m(K)$ here turns out to be the Chandrasekhar mass for the model. If the mass is too large, gravitational collapse may happen.
Notice Lemma 2.2 also needs to be modified if we assume (F3’) instead of (F3). Then in the proof of Lemma 2.6, a potential issue with a large value of mass is that it may invalidate the earlier assertion that $\sigma_{m}{}^{\frac{1}{3}}\leq cV_{\sigma_{m}}$ a.e. on the set where $\sigma_{m}\geq K$ for some $K$ large enough”. It may also cause issues in the proof of Lemma 2.22 below, see also Remark 2.24. For more discussion, one can see [16, Section 2] or [1, Section 1,6,8 and Appendix].
@@
Currently, we do not know whether the minimizer $\sigma_{m}$ with fixed mass $m$ is unique. Further discussion regarding uniqueness (up to translation) will be conducted under additional assumption later. However, even if multiple minimizers $\sigma_{m}^{1},\sigma_{m}^{2}\ldots\sigma_{m}^{i}\ldots$ exist under the present assumptions, due to Theorem 2.6 (iv), together with interpolation inequality Proposition A.2, we know that all of them have uniform $L^{p}$ bound. Moreover, according to Theorem 2.6 (viii) and Remark 2.13, we observe their corresponding chemical potential $\lambda_{m}^{i}$ are also uniformly bounded by $e^{\prime}_{0}(m^{-})$, which is strictly negative. This observation is applied in [16, Section 6] to estimate the size of rotating stars by showing that they are close to non‑rotating stars when their mutual distances are large—a conclusion established through an energy convergence.
Remark 2.15.
Notice Auchmuty and Beals only concluded that (EP’) is satisfied in the region where $\rho>0$ [1, Theorem A], whereas here we have established a stronger result—namely, that the equation holds over the entire $\mathbb{R}^{3}$ space.
Remark 2.16.
We point out that $\sigma_{m}$ is a minimizer on $m{R}\left(\mathbb{R}^{3}\right)$ is important to show the global continuity of $\sigma_{m}$, i.e. $\sigma_{m}$ is continuous in $\mathbb{R}^{3}$. Because there is no constraint on the support of perturbations, the set of admissible perturbations here can be taken exactly as $P_{\infty}\left(\sigma_{m}\right)$ instead of something like $P_{loc}$ in [7, Subection 5.2]. It means $E_{0}^{\prime}\left(\sigma_{m}\right)\geq\lambda_{m}$ a.e. in $\mathbb{R}^{3}$ and then $\sigma_{m}=\phi\circ\left[V_{\sigma_{m}}(x)+\lambda_{m}\right]_{+}$ a.e. in $\mathbb{R}^{3}$ which implies the global continuity.
However, if $\sigma$ is a “constrained” minimizer such as $\sigma_{R}$ mentioned above, or $\rho(m)$ given in [6, Theorem 3.1.15], we might only know the continuity of $\rho$ in $\{\rho>0\}$ or in the constrained domain, and whether $\nabla P(\rho)$ exists on the boundary $\partial\{\rho>0\}$ is not clear. One way to show the global continuity is to show such “constrained” minimizer can indeed be a “global” minimizer, for example we will show in single star case that $\sigma_{R}$ can be a minimizer on $m{R}\left(\mathbb{R}^{3}\right)$ in the following when $R$ large enough. However, when it comes to binary star system or star-planet system, the global minimal energy cannot be attained as in Remark 2.45 and Remark 2.46 in the following, or [7, Section 2]. It turns out once we know $\sigma$ is not only a constrained minimizer but also a Wasserstain $L^{\infty}$ local minimizer, we can still show the global continuity as [7, Theorem 2.25 (iii)], which shows the power of $W^{\infty}$ metric. More discussion about how to show $\sigma$ is a Wasserstein $L^{\infty}$ local minimizer can be seen in [6].
Before showing the existence of minimizer on $m{R}\left(\mathbb{R}^{3}\right)$, it is worth noting the fact that if we replace $\frac{4}{3}$ by any smaller number in (F3), there is no minimizer since the energy may be unbounded below. The idea comes from [1, Remark in Section 8].
@@
Follow the proof of L’Hôpital’s rule [20, 5.13 Theorem], we know $\lim\limits_{\rho\rightarrow\infty}\sup P(\rho)\rho^{-\gamma}=K$ implies $\lim\limits_{\rho\rightarrow\infty}\sup A(\rho)\rho^{-\gamma}\leq 3K$. then we know $\lim\limits_{\rho\rightarrow\infty}A(\rho)\rho^{-\frac{4}{3}}=$ 0. Therefore, $\forall\hat{\epsilon}>0,\exists S_{\hat{\epsilon}}>0$, such that $A(s)\leq\hat{\epsilon}s^{-\frac{4}{3}}$ when $s\geq S_{\hat{\epsilon}}$. In the statement of Proposition A.4, let $\hat{c}$ be the smallest constant such that $\left|\int\rho V_{\rho}\right|\leq 2\hat{c}\left(\int|\rho|\right)^{\frac{2}{3}}\int|\rho|^{\frac{4}{3}}$ for all $\rho\in L^{1}\cap L^{\frac{4}{3}}$, then $\hat{c}>0$. Otherwise $\left|\int\rho V_{\rho}\right|\leq 0$ implies $\rho=0$, but $\{0\}$ is strictly contained in $L^{1}\cap L^{\frac{4}{3}}$.
@@
$U\left(\rho_{\delta}\right)$ $=\int A\left(\rho_{\delta}\right)$
$=\int_{\left\{\rho_{\delta}\geq S_{\hat{\epsilon}}\right\}}A\left(\rho_{\delta}\right)+\int_{\left\{\rho_{\delta}<S_{\hat{\epsilon}\}}\right.}A\left(\rho_{\delta}\right)$
@@
Notice that given $\hat{\epsilon}$ and $\epsilon,\mu\left(\left\{\rho_{\delta}<S_{\hat{\epsilon}}\right\}\right)=\mu\left(\left\{x\left\lvert\,\rho_{\epsilon}\left(\delta^{-1}x\right)<\frac{S_{\hat{\epsilon}}\delta^{3}}{m}\right.\right\}\right)=\delta^{3}\mu\left(y\left\lvert\,\left\{\rho_{\epsilon}(y)<\frac{S_{\hat{\epsilon}}\delta^{3}}{m}\right\}\right.\right)\rightarrow 0$ as $\delta\rightarrow 0$. Then we pick $\epsilon=\frac{\hat{c}}{2},\hat{\epsilon}=\frac{m^{\frac{2}{3}}\hat{c}}{4}$, and then when $\delta\rightarrow 0$
@@
$\leq\frac{m^{2}\hat{c}}{4\delta}\int_{\mathbb{R}^{3}}\rho_{\epsilon}^{\frac{4}{3}}+A\left(S_{\hat{\epsilon}}\right)\mu\left(\left\{\rho_{\delta}<S_{\hat{\epsilon}}\right\}\right)-\frac{m^{2}\hat{c}}{2\delta}\int_{\mathbb{R}^{3}}\rho_{\epsilon}^{\frac{4}{3}}\rightarrow$-$\infty$
@@
Actually, there is another way to see the unboundedness. From $\lim\limits_{\rho\rightarrow\infty}\sup A(\rho)\rho^{-\gamma}\leq 3K$ we can notice $A(\rho)\leq(3K+\hat{\epsilon})\rho^{\gamma}$ for large $\rho$, then similarly as above we have $E_{0}\left(\rho_{\delta}\right)=U\left(\rho_{\delta}\right)-\frac{G\left(\rho_{\delta},\rho_{\delta}\right)}{2}\leq m^{\gamma}\delta^{-3+3\gamma}(3K+\hat{\epsilon})\int\rho_{\epsilon}^{\gamma}+A\left(S_{\hat{\epsilon}}\right)\mu\left(\left\{\rho_{\delta}<S_{\hat{\epsilon}}\right\}\right)-\frac{(\hat{c}-\epsilon)m^{2}}{\delta}\int\left|\rho_{\epsilon}\right|^{\frac{4}{3}}$. Notice $3+3\gamma<-1$, thus we can choose suitable $\epsilon$ and $\hat{\epsilon}$ such that $E_{0}\left(\rho_{\delta}\right)\rightarrow$-$\infty$ as $\delta\rightarrow 0$. From this point we can know that the number $\gamma=\frac{4}{3}$ in (F3) is chosen such that $3+3\gamma\geq-1$, then (F3) can help the blowing up of $U(\rho)$ to control the gravitational collapse $G(\rho,\rho)$, as discussion in [7, Remark 2.2].
In fact, in the case that blowing up of $U(\rho)$ is too large, with the same strategy we can show the energy may be also unbounded above. That is, if $\frac{4}{3}$ is replaced by any larger number in (F3), then we will have $U\left(\rho_{\delta}\right)>>G\left(\rho_{\delta},\rho_{\delta}\right)>>1$.
@@
In fact, the energy may be also unbounded below or above even if we replace ${R}\left(\mathbb{R}^{3}\right)$ by ${R}_{p}\left(\mathbb{R}^{3}\right):=\left\{\rho\in L^{p}\left(\mathbb{R}^{3}\right)\mid\rho\geq 0,\int\rho=1\right\}$ for some $p\leq 0$, since $C_{c}^{0}\left(\mathbb{R}^{3}\right)\subset L^{p}$ for any $p\geq 1$, and then $\rho_{\epsilon}\in{R}_{p}\left(\mathbb{R}^{3}\right)$.
@@
If (F3) is replaced by the assumption $\lim\limits_{\rho\rightarrow\infty}\inf P(\rho)\rho^{-\frac{4}{3}}=K>0$, Auchmuty and Beals [1, Section 1 and 6] show there is a constant $M_{0}>0$ such that $\forall 0\leq m<M_{0}$ there is a minimizer on $m{R}\left(\mathbb{R}^{3}\right)$. If $\lim\limits_{\rho\rightarrow\infty}\inf P(\rho)\rho^{-\frac{4}{3}}=K>0$ is strengthened to $\lim\limits_{\rho\rightarrow\infty}P(\rho)\rho^{-\frac{4}{3}}=K>0$, then $\forall m>M_{0}$, the energy is unbounded below on $m{R}\left(\mathbb{R}^{3}\right)$ [1, Remark in Section 8]. Moreover, they [1, Remark in Section 6, and Appendix] also show that if $P$ is the function occurring in the theory of white dwarf stars (relativistically degenerate gases) $M_{0}$ turns out to be the same as the limiting mass $M_{3}$ in Chandrasekhar’s theory [4].
@@
Assume $\sigma_{R}$ is the energy minimizer in $W_{R}$, which is given in (17). Then the $L^{\infty}\left(\mathbb{R}^{3}\right)$ bound of $\sigma_{R}$ is uniform, i.e. $\exists C>0$, such that $\forall R>0$, $\left\|\sigma_{R}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}\leq C$.

### v2

@@ 2.2 Existence Results for Non-rotating Bodies
For Part (vii), we already know $\lambda_{m}<0$. Moreover, (17) (18), together with $\sigma_{m}$ is continuous, give (EL), see for example [46, Section 5] or [17, Section 2]. (Part (vii))
@@
For Part (ii), we only consider the case that complies with polytropic equations of state $P(\sigma)=K\sigma^{\gamma}$. The general case can be found in [42, Theorem 3 (d)] [^4] However, it should be noted that the argument there is set within a quantum mechanical framework and does not include an internal energy term, although McCann mentioned that the results can be applied “equally well to all $A(\rho)$ consistent with (F1)-(F3)[46, Section 3].. We utilize the scaling relations mentioned in Section 3. That is, given $\sigma$ a non-rotating minimizer with mass $1$, then $\sigma_{m}(x)=\frac{1}{A}\sigma\left(\frac{1}{B}x\right)$ will be a non-rotating minimizer with mass $m$, where $A(m)=m^{-\frac{2}{3\gamma-4}}$, $B(m)=m^{\frac{\gamma-2}{3\gamma-4}}$. At the same time, we know $e_{0}(m)=E_{0}\left(\sigma_{m}\right)=E_{0}\left(\frac{1}{A}\sigma\left(\frac{1}{B}x\right)\right)=m^{\frac{5\gamma-6}{3\gamma-4}}e_{0}$. Notice when $m=0$, such expression also holds since $e_{0}(0)=0$. By the similar arguments as the proof of Step (2) (Lemma 2.24), we know $e_{0}\coloneq e_{0}(1)<0$, which implies $e_{0}(m)$ decreases continuously from $e_{0}(0)=0$ and is strictly concave (Part (ii)).
@@
For Part (viii), first the concavity of $e_{0}(m)$ implies the existence of left-hand and right-hand derivations (since they are monotonic), and they are equal a.e. (since concavity implies Lipschitz continuity and then apply Rademacher’s theorem [21, Subsection 5.8.3]). Thanks to Euler-Lagrange equations (EL), we multiply $\sigma_{m}$ on both sides and integrate, and then we obtain
$\int_{\mathbb{R}^{3}}E_{0}^{\prime}\left(\sigma_{m}\right)\sigma_{m}\,dx=\lambda_{m}\cdot m$
Notice due to Part (iv) and (v), we know $\sigma_{m}\in P_{\infty}\left(\sigma_{m}\right)$ thus $\int_{\mathbb{R}^{3}}E_{0}^{\prime}\left(\sigma_{m}\right)\sigma_{m}\,dx$ makes sense and
$\int_{\mathbb{R}^{3}}E_{0}^{\prime}\left(\sigma_{m}\right)\sigma_{m}\,dx=\lim\limits_{t\rightarrow 0^{+}}\frac{E_{0}\left(\sigma_{m}+t\sigma_{m}\right)-E_{0}\left(\sigma_{m}\right)}{t}=\lim\limits_{t\rightarrow 0^{-}}\frac{E_{0}\left(\sigma_{m}\right)-E_{0}\left(\sigma_{m}-t\sigma_{m}\right)}{t}$
Since $\sigma_{m}+t\sigma_{m}$ has mass $(1+t)m$, then
$e_{0}((1+t)m)=E_{0}\left(\sigma_{(1+t)m}\right)\leq E_{0}\left(\sigma_{m}+t\sigma_{m}\right)$
Thus
$\lim\limits_{t\rightarrow 0^{+}}\frac{E_{0}\left(\sigma_{m}+t\sigma_{m}\right)-E_{0}\left(\sigma_{m}\right)}{t}\geq\lim\limits_{t\rightarrow 0^{+}}\frac{e_{0}((1+t)m)-e_{0}(m)}{t}=me_{0}^{\prime}\left(m^{+}\right)$
Hence we have $\lambda_{m}\geq e_{0}^{\prime}\left(m^{+}\right)$. Similarly, we have $\lambda_{m}\leq e_{0}^{\prime}\left(m^{-}\right)$. (Part (viii))
@@
For Part (ix), thanks to Part (vi), we know the global continuity of $\sigma_{m}$, and then we can apply the same arguments as in [17, Section 2] and show the existence of $\nabla P\left(\sigma_{m}\right)$ on the boundary $\partial\left\{\sigma_{m}>0\right\}$, and verify $\sigma_{m}$ is a solution to reduced Euler-Poisson equations (EP’) in $\mathbb{R}^{3}$. (Part (ix)) ∎
@@
When we prove Theorem 2.6 (vi), we mention $\sigma_{m}{}^{\frac{1}{3}}\leq cV_{\sigma_{m}}$ a.e. on the set where $\sigma_{m}\geq K$ for some $K$ large enough. Actually it turns out even if we do not know $\lambda_{m}<0$, $\sigma_{m}{}^{\frac{1}{3}}\leq cV_{\sigma_{m}}$ can still hold true since we can absorb $\lambda_{m}$ when $\sigma_{m}$ is large enough. See [4, Lemma 3].
@@
We point out in Theorem 2.6 we can show the existence of non-rotating single star no matter how large the mass is. However, if we replace (F3) by (F3’), then we can only obtain minimizers with total mass not larger than $m(K)$ via variational method for some $m(K)>0$. $m(K)$ here turns out to be the Chandrasekhar mass for the model. If the mass is too large, gravitational collapse may happen [4, Appendix].
To clarify, replacing (F3) with (F3’) would require more careful treatment to prove Lemma 2.6. First, observe that Lemma 2.2 needs to be modified under assumption (F3’) rather than (F3). In the proof of Lemma 2.6, a potential issue with a large mass is that it may invalidate the earlier assertion that $\sigma_{m}{}^{\frac{1}{3}}\leq cV_{\sigma_{m}}$ a.e. on the set where $\sigma_{m}\geq K$ for some $K$ large enough”. Moreover, (F3’) may also lead to a more delicate treatment in the proof of Lemma 2.21 below; see also Remark 2.23. For more discussion, one can see [46, Section 2] or [4, Section 1,6,8 and Appendix].
@@
Currently, we do not know whether the minimizer $\sigma_{m}$ with fixed mass $m$ is unique. Further discussion regarding uniqueness (up to translation) will be conducted under additional assumptions later. However, even if multiple minimizers $\sigma_{m}^{1},\sigma_{m}^{2}\ldots\sigma_{m}^{i}\ldots$ exist under the present assumptions, due to Theorem 2.6 (iv), together with interpolation inequality [8, Section 4.2], we know that all of them have uniform $L^{p}$ bound. Moreover, according to Theorem 2.6 (viii) and Remark 2.14, we observe their corresponding Lagrange multipliers $\lambda_{m}^{i}$ are also uniformly bounded by $e^{\prime}_{0}(m^{-})$, which is strictly negative. This observation is applied in [46, Section 6].
Before showing the existence of a minimizer on $m{R}\left(\mathbb{R}^{3}\right)$, it is worth noting that if we replace $\frac{4}{3}$ by any smaller number in (F3), there is no minimizer since the energy may be unbounded below. The idea comes from [4, Remark in Section 8].
@@
Follow the proof of L’Hôpital’s rule [52, Theorem 5.13], we know $\lim\limits_{\rho\rightarrow\infty}\sup P(\rho)\rho^{-\gamma}=K$ implies $\lim\limits_{\rho\rightarrow\infty}\sup A(\rho)\rho^{-\gamma}\leq 3K$. then we know $\lim\limits_{\rho\rightarrow\infty}A(\rho)\rho^{-\frac{4}{3}}=$ 0. Therefore, $\forall\hat{\epsilon}>0,\exists S_{\hat{\epsilon}}>0$, such that $A(s)\leq\hat{\epsilon}s^{-\frac{4}{3}}$ when $s\geq S_{\hat{\epsilon}}$. In the statement of Proposition A.2, let $\hat{c}$ be the smallest constant such that $\forall\rho\in L^{1}\cap L^{\frac{4}{3}}$, we have
$\left|\int_{\mathbb{R}^{3}}\rho V_{\rho}\,dx\right|\leq 2\hat{c}\left(\int_{\mathbb{R}^{3}}|\rho|\,dx\right)^{\frac{2}{3}}\int_{\mathbb{R}^{3}}|\rho|^{\frac{4}{3}}\,dx$
Then $\hat{c}>0$. Otherwise $\left|\int_{\mathbb{R}^{3}}\rho V_{\rho}\,dx\right|\leq 0$ implies $\rho=0$, but 0 is not the only element in $L^{1}\cap L^{\frac{4}{3}}$.
@@
$\left|\int_{\mathbb{R}^{3}}\widetilde{\rho_{\epsilon}}V_{\widetilde{\rho_{\epsilon}}}\,dx\right|>2\left(\hat{c}-\frac{\epsilon}{2}\right)\left(\int_{\mathbb{R}^{3}}\left|\widetilde{\rho_{\epsilon}}\right|\,dx\right)^{\frac{2}{3}}\int_{\mathbb{R}^{3}}\left|\widetilde{\rho_{\epsilon}}\right|^{\frac{4}{3}}\,dx$
@@
$\frac{\left|\int_{\mathbb{R}^{3}}\rho_{\epsilon}V_{\rho_{\epsilon}}\,dx\right|}{\left(\int_{\mathbb{R}^{3}}\left|\rho_{\epsilon}\right|\,dx\right)^{\frac{2}{3}}\int_{\mathbb{R}^{3}}\left|\rho_{\epsilon}\right|^{\frac{4}{3}}\,dx}>2(\hat{c}-\epsilon)$
We can also write it as
$G\left(\rho_{\epsilon},\rho_{\epsilon}\right)=\left|\int_{\mathbb{R}^{3}}\rho_{\epsilon}V_{\rho_{\epsilon}}\,dx\right|>2(\hat{c}-\epsilon)\left(\int_{\mathbb{R}^{3}}\left|\rho_{\epsilon}\right|\,dx\right)^{\frac{2}{3}}\int_{\mathbb{R}^{3}}\left|\rho_{\epsilon}\right|^{\frac{4}{3}}\,dx$
It implies $\rho_{\epsilon}\neq 0$. We can also assume $\rho_{\epsilon}\geq 0$ since we can replace $\rho_{\epsilon}$ by $\left[\rho_{\epsilon}\right]_{+}$ if needed. Without loss of generality, we assume $\int_{\mathbb{R}^{3}}\rho_{\epsilon}\,dx=1$. Let $\rho_{\delta}(x)=m\delta^{-3}\rho_{\epsilon}\left(\delta^{-1}x\right)$, $\delta>0$. Then $\rho_{\delta}\in m{R}\left(\mathbb{R}^{3}\right)$ since $C_{c}^{0}\left(\mathbb{R}^{3}\right)\subset L^{\frac{4}{3}}$. Similar as discussions in Section 3, we can show $G\left(\rho_{\delta},\rho_{\delta}\right)=\frac{m^{2}}{\delta}G\left(\rho_{\epsilon},\rho_{\epsilon}\right)$, and
$U\left(\rho_{\delta}\right)$ $=\int_{\mathbb{R}^{3}}A\left(\rho_{\delta}\right)\,dx$
$=\int_{\left\{\rho_{\delta}\geq S_{\hat{\epsilon}}\right\}}A\left(\rho_{\delta}\right)\,dx+\int_{\left\{\rho_{\delta}<S_{\hat{\epsilon}\}}\right.}A\left(\rho_{\delta}\right)\,dx$
$\leq\int_{\left\{\rho_{\delta}\geq S_{\hat{\epsilon}}\right\}}\hat{\epsilon}\cdot\rho_{\delta}^{\frac{4}{3}}\,dx+\int_{\left\{\rho_{\delta}<S_{\hat{\epsilon}\}}\right.}A\left(S_{\hat{\epsilon}}\right)\,dx$
@@
Here $\mu$ denotes the Lebesgue measure. Notice that given $\hat{\epsilon}$ and $\epsilon$, as $\delta\rightarrow 0$, we have
$\mu\left(\left\{\rho_{\delta}<S_{\hat{\epsilon}}\right\}\right)=\mu\left(\left\{x\left\lvert\,\rho_{\epsilon}\left(\delta^{-1}x\right)<\frac{S_{\hat{\epsilon}}\delta^{3}}{m}\right.\right\}\right)=\delta^{3}\mu\left(y\left\lvert\,\left\{\rho_{\epsilon}(y)<\frac{S_{\hat{\epsilon}}\delta^{3}}{m}\right\}\right.\right)\rightarrow 0$
Then we pick $\epsilon=\frac{\hat{c}}{2},\hat{\epsilon}=\frac{m^{\frac{2}{3}}\hat{c}}{4}$, and then when $\delta\rightarrow 0$
@@
$\leq\frac{m^{2}\hat{c}}{4\delta}\int_{\mathbb{R}^{3}}\rho_{\epsilon}^{\frac{4}{3}}\,dx+A\left(S_{\hat{\epsilon}}\right)\mu\left(\left\{\rho_{\delta}<S_{\hat{\epsilon}}\right\}\right)-\frac{m^{2}\hat{c}}{2\delta}\int_{\mathbb{R}^{3}}\rho_{\epsilon}^{\frac{4}{3}}\,dx\rightarrow$-$\infty$
@@
Actually, there is another way to see the unboundedness. From $\lim\limits_{\rho\rightarrow\infty}\sup A(\rho)\rho^{-\gamma}\leq 3K$, we can notice $A(\rho)\leq(3K+\hat{\epsilon})\rho^{\gamma}$ for large $\rho$, then similarly as above we have
$\leq m^{\gamma}\delta^{-3+3\gamma}(3K+\hat{\epsilon})\int_{\mathbb{R}^{3}}\rho_{\epsilon}^{\gamma}\,dx+A\left(S_{\hat{\epsilon}}\right)\mu\left(\left\{\rho_{\delta}<S_{\hat{\epsilon}}\right\}\right)-\frac{(\hat{c}-\epsilon)m^{2}}{\delta}\int_{\mathbb{R}^{3}}\left|\rho_{\epsilon}\right|^{\frac{4}{3}}\,dx$
Notice $3+3\gamma<-1$, thus we can choose suitable $\epsilon$ and $\hat{\epsilon}$ such that $E_{0}\left(\rho_{\delta}\right)\rightarrow$-$\infty$ as $\delta\rightarrow 0$. From this point we can know that the number $\gamma=\frac{4}{3}$ in (F3) is chosen such that $3+3\gamma\geq-1$, then (F3) can help the blowing up of $U(\rho)$ to control the gravitational collapse $G(\rho,\rho)$.
In fact, in the case that blowing up of $U(\rho)$ is too large, with the same strategy we can show the energy may also be unbounded above. That is, if $\frac{4}{3}$ is replaced by any larger number in (F3), then we will have $U\left(\rho_{\delta}\right)\gg G\left(\rho_{\delta},\rho_{\delta}\right)\gg 1$.
@@
Thanks to the arguments above, we can first find a $\rho\in C_{c}^{0}\left(\mathbb{R}^{3}\right)$ with mass 1, and set $\rho_{\delta}(x)=m\delta^{-3}\rho\left(\delta^{-1}x\right)$. If $K<\infty$, notice
$\left\{x\,\middle|\,\rho_{\delta}\geq S_{\hat{\epsilon}}\right\}=\left\{x\,\middle|\,\rho\left(\delta^{-1}x\right)\geq\frac{S_{\hat{e}}\delta^{3}}{m}\right\}$
by a change of variable, we know for $\delta$ large enough, we have
$U\left(\rho_{\delta}\right)\geq m^{\gamma}\delta^{-3+3\gamma}(3K-\hat{\epsilon})\int_{\left\{\rho\geq\frac{S_{\hat{\epsilon}}\delta^{3}}{m}\right\}}\rho^{\gamma}\,dx\geq m^{\gamma}\delta^{-3+3\gamma}(3K-\hat{\epsilon})\int_{\{\rho\geq C\}}\rho^{\gamma}\,dx$
Notice we can find a $C$ such that $\int_{\{\rho\geq C\}}\rho^{\gamma}\,dx>0$ due to the same arguments in [17, Section 4]. By Proposition A.2, together with $\gamma>\frac{4}{3}$, we know as $\delta\rightarrow 0$,
$E_{0}\left(\rho_{\delta}\right)=U\left(\rho_{\delta}\right)-\frac{G\left(\rho_{\delta},\rho_{\delta}\right)}{2}\geq m^{\gamma}\delta^{-3+3\gamma}(3K-\hat{\epsilon})\int_{\{\rho\geq C\}}\rho^{\gamma}\,dx-\frac{\hat{c}m^{2}}{\delta}\int_{\mathbb{R}^{3}}|\rho|^{\frac{4}{3}}\,dx\rightarrow+\infty$
@@
In fact, the energy may be also unbounded below or above even if we replace ${R}\left(\mathbb{R}^{3}\right)$ by ${R}_{p}\left(\mathbb{R}^{3}\right):=\left\{\rho\in L^{p}\left(\mathbb{R}^{3}\right)\mid\rho\geq 0,\int_{\mathbb{R}^{3}}\rho\,dx=1\right\}$ for some $p\geq 1$, since $C_{c}^{0}\left(\mathbb{R}^{3}\right)\subset L^{p}$ for any $p\geq 1$, and then $\rho_{\epsilon}\in{R}_{p}\left(\mathbb{R}^{3}\right)$.
@@
If (F3) is replaced by the assumption $\lim\limits_{\rho\rightarrow\infty}\inf P(\rho)\rho^{-\frac{4}{3}}=K>0$, Auchmuty and Beals [4, Section 1 and 6] show there is a constant $M_{0}>0$ such that $\forall 0\leq m<M_{0}$ there is a minimizer on $m{R}\left(\mathbb{R}^{3}\right)$. If $\lim\limits_{\rho\rightarrow\infty}\inf P(\rho)\rho^{-\frac{4}{3}}=K>0$ is strengthened to $\lim\limits_{\rho\rightarrow\infty}P(\rho)\rho^{-\frac{4}{3}}=K>0$, then $\forall m>M_{0}$, the energy is unbounded below on $m{R}\left(\mathbb{R}^{3}\right)$ [4, Remark in Section 8]. Moreover, they [4, Remark in Section 6, and Appendix] also show that if $P$ is the function occurring in the theory of white dwarf stars (relativistically degenerate gases), $M_{0}$ turns out to be the same as the limiting mass $M_{3}$ in Chandrasekhar’s theory [12].
@@
Assume $\sigma_{R}$ is the energy minimizer in $W_{R}$, which is given in (16). Then the $L^{\infty}\left(\mathbb{R}^{3}\right)$ bound of $\sigma_{R}$ is uniform, i.e., $\exists C>0$, such that $\forall R>0$, $\left\|\sigma_{R}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}\leq C$.